two consecutive odd intergers have a product of 99. find the larger number.
Answer by nerdybill(7384)

two consecutive odd intergers have a product of 99. find the larger number.
.
Let x = smaller consecutive odd integer
then
x+2 = larger consecutive odd integer
.
x(x+2) = 99
x^2 + 2x = 99
x^2 + 2x - 99 = 0
(x+11)(x-9) = 0
x = {-11,9}
.
the larger odd integers would then be:
x = {-9,11}
.
Therefore, TWO possible sets:
-11 and -9
OR
9 and 11
.
Therefore, the answer could either be -9 or 11
If they want the positive solution it would be 11